OpenStudy (anonymous):

Help!!! will give medal Multiply. x+2/5x-15 times (10x^2-25x-15)

OpenStudy (anonymous):

Choices A. (2x+1)/5(x-3) B. (2x+1)(x-3)/5 C.(10x+5)(x+2) D. (x+2)(2x+1)

OpenStudy (campbell_st):

well factor the quadratic equation and you get \[5(2x^2 -5x - 3)\] which can be further factored to \[5(x - 3)(2x + 1)\] so this can be written as \[(5x - 15)(2x + 1)\] so the problem is now \[\frac{(x + 2)}{(5x - 15)} \times \frac{(5x - 15)(2x + 1)}{1}\] cancel common factors and you'll have your solution..

OpenStudy (anonymous):

what are those @campbell_st

OpenStudy (campbell_st):

well what do you see that is the same... in the numerator and denominator.... whatever it is... cross it out... what's left is the answer

OpenStudy (anonymous):

5x is the same so I cross out tht

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):